Hastings Co. sells widgets. The company is currently developing a purchases budget for July. Projected sales volumes are 154How much cash outflow for purchases will Hastings budget for July? (Round to nearest dollar if not a whole number.) Answer:
Answer #1

Answer:

Budgeted for June = 154

For July = 183

For August = 195

As per term company purchase 21 % of the inventory for the next month's sale volume.

Purchase in June = 154 + 38 approx(21 % of 183)

Purchase in July = (183 -38  as already purchased in june) + 41 approx ( 21 % 195)

= 145 + 41

= 186

1. Purchase in July = 186 widgets

2. Total purhcase = 3906 (186*21)

3. Cash outflow = 3906

Hit like if it was helpful